# A Parameter-Efficient Tuning Framework for Language-guided Object Grounding and Robot Grasping

### v1

@@ Abstract
The language-guided robot grasping task requires a robot agent to integrate multimodal information from both visual and linguistic inputs to predict actions for target-driven grasping. While recent approaches utilizing Multimodal Large Language Models (MLLMs) have shown promising results, their extensive computation and data demands limit the feasibility of local deployment and customization. To address this, we propose a novel CLIP-based [1] multimodal parameter-efficient tuning (PET) framework designed for three language-guided object grounding and grasping tasks: (1) Referring Expression Segmentation (RES), (2) Referring Grasp Synthesis (RGS), and (3) Referring Grasp Affordance (RGA). Our approach introduces two key innovations: a bi-directional vision-language adapter that aligns multimodal inputs for pixel-level language understanding and a depth fusion branch that incorporates geometric cues to facilitate robot grasping predictions. Experiment results demonstrate superior performance in the RES object grounding task compared with existing CLIP-based full-model tuning or PET approaches. In the RGS and RGA tasks, our model not only effectively interprets object attributes based on simple language descriptions but also shows strong potential for comprehending complex spatial reasoning scenarios, such as multiple identical objects present in the workspace.
@@ I Introduction
Recent robot grasping studies using Large Language Models (LLMs) or Large Multimodal Language Models (MLLMs) have demonstrated strong object grounding and grasp detection performance [16, 17, 18, 19, 20]. However, full-model training or fine-tuning is computationally expensive, limiting the feasibility of local deployment. In this paper, we propose a parameter-efficient tuning (PET) approach [21, 22] for object grounding and robot grasping, enabling competitive performance by incorporate a lightweight module into a frozen pre-trained model.
@@ IV Experiments
TABLE I: Performance comparison of different SOTA CLIP-based models on RefCOCO related datasets using mIoU metric. Backbone: visual encoder architecture. U: the UMD test split for RefCOCO(g) dataset. Params: total number of parameters to be tuned in the corresponding baselines.
@@
ETOG (Ours) CLIP-R50 72.31 75.49 66.62 63.26 69.46 53.27 60.80 61.98 65.40 32.93M
@@
ETOG (Ours) CLIP-R101 73.37 76.16 68.54 64.70 70.86 54.94 63.06 63.66 66.91 30.57M
@@
ETOG (Ours) CLIP-ViT-B 73.37 76.90 69.34 65.95 71.47 56.94 63.75 64.63 67.79 29.89M
@@
ETOG (Ours) ResNet50 70.11 73.97 63.76 58.63 65.66 47.86 57.35 59.87 62.15
ETOG (Ours) ResNet101 71.11 74.30 65.61 60.00 67.27 49.73 60.09 61.23 63.67
ETOG (Ours) ViT-B 71.35 76.14 66.73 62.33 68.52 51.90 61.12 62.85 65.12
@@ IV-A Referring Expression Segmentation for Object Grounding
ETOG Comparison with RES Baselines. In Table II, we compare our ETOG with the state-of-the-art CLIP-based RES methods on RefCOCO datasets using mIoU metric. ETOG outperforms all baseline methods across different backbone architectures. CLIPORT-semantic, a variant of CLIPORT [30], excludes the depth processing branch and uses simple element-wise multiplication to fuse vision and language features. The comparison between CLIPORT-semantic and our method highlights the effectiveness of our VL-adapter. In particular, with the CLIP [1] ResNet-101 [37] backbone, our method–using only $1.9$M tunable parameters–achieves a $2.66\%$ and $1.95\%$ performance gain across 8 evaluation tasks compared to CRIS [25] and EAVL [26] that are fully CLIP fine-tuned methods. Additionally, ETOG with $1.21$M tunable ViT-B-16 backbone parameters shows a $4.35\%$ and $0.73\%$ performance improvement compared to the latest parameter-efficient tuning approaches, ETRIS [21] and BARLERIA [22], respectively. Table II further compares our model with full fine-tuned RES models, showing that ETOG achieves competitive object grounding performance even against non-CLIP-based approaches. Fig. 3 provides qualitative results, including model predictions and feature map visualizations, with the fourth column showting the attention regions of the well-fused multimodal feature $f_{c}$.
@@ IV-B Referring Grasp Synthesis with Depth Fusion
Ablation Study on Depth Fusion Branch. One primary SOTA baseline we evaluate against is CROG [11], a CLIP-based RGS method with full model fine-tuned with OCID-VLG dataset. As shown in Table IV, ETRG-A without depth model (ResNet-50) achieves a slightly better ($+3.96\%$ $J@1$) performance and competitive object grounding (RES) performance compared CROG. After integrating depth inputs into VL-adapter (details in Fig. 2) fusing the object geometric information into visual-language branches, our ETRG-A with depth fusion full model attains higher grasping rectangle prediction ($+11.86\%$ $J@1$) accuracy with ResNet-50 backbone and ($+12.18\%$ $J@1$) accuracy with ResNet-101 backbone compared with CROG. CROG is marginally better than our models in the RES task, but we attribute this to the simplicity of the OCID-VLG dataset, which primarily features tabletop objects and lacks the scene diversity found in RefCOCO-related datasets. This environment favors full model fine-tuning methods like CROG, potentially leading to overfitting. Although CROG has 147M tunable parameters which is 3 times larger than our proposed depth-fusion approach, our ETRA-A model still achieves competitive performance on this RES task.
@@ IV-C Referring Grasp Affordance for Spatial Reasoning
ETRG-B (Ours) 90.75 86.56 88.38 88.56
Simulation Results. We evaluate our methods on affordance prediction with spatial reasoning. The experiments include two primary scene types: Absolute and Relative. In each test scene, there are three identical object instances, and spatial language expression are used to specify the target. To increase the difficulty of spatial reasoning, we create test cases with four objects and seven objects. As illustrated in Table V, both the proposed methods, with and without depth fusion, demonstrate strong performance in handling complex spatial reasoning tasks. Our model with depth fusion achieves slightly higher performance compared to the RGB-only method.
@@
Real Robot Results. We conducted experiments using a Franka Emika Panda robot equipped with a FESTO DHAS soft gripper. We collected over 10 household objects and created scenes with 6 randomly selected objects. Each scene included up to 3 duplicated instances of one category to test spatial reasoning. For each baseline, we report object grounding accuracy, which assesses whether the grasp is on the correct target, and the overall grasp success rate. A detailed visualization of correct mustard bottle identification with feasible grasp poses is shown in Fig. 4. Across 40 grasp attempts with different object arrangements (Table VI), both methods achieve comparable grounding performance, with our model incorporating depth information showing slightly better grasping accuracy.
TABLE VI: ETRG-B Real robot results on RGA spatial reasoning. Evaluated with grounding accuracy and grasp success rate.
@@ V Conclusion
In this work, we presented a novel CLIP-based parameter-efficient tuning framework that can be adapted to a range of tasks, including RES, RGS, and RGA tasks. Our method incorporated a vision-language fusion adapter and a depth-fusion branch, enabling effective open-vocabulary object grounding and robot grasping. We further explored the spatial reasoning capabilities in robot grasping, which the original CLIP [1] model does not address. The proposed models demonstrated competitive performance across these challenging vision language and robotics tasks.

### v3

@@ Abstract
The language-guided robot grasping task requires a robot agent to integrate multimodal information from both visual and linguistic inputs to predict actions for target-driven grasping. While recent approaches utilizing Multimodal Large Language Models (MLLMs) have shown promising results, their extensive computation and data demands limit the feasibility of local deployment and customization. To address this, we propose a novel CLIP-based [1] multimodal parameter-efficient tuning (PET) framework designed for three language-guided object grounding and grasping tasks: (1) Referring Expression Segmentation (RES), (2) Referring Grasp Synthesis (RGS), and (3) Referring Grasp Affordance (RGA). Our approach introduces two key innovations: a bi-directional vision-language adapter that aligns multimodal inputs for pixel-level language understanding and a depth fusion branch that incorporates geometric cues to facilitate robot grasping predictions. Experiment results demonstrate superior performance in the RES object grounding task compared with existing CLIP-based full-model tuning or PET approaches. In the RGS and RGA tasks, our model not only effectively interprets object attributes based on simple language descriptions but also shows strong potential for comprehending complex spatial reasoning scenarios, such as multiple identical objects present in the workspace. Project page: https://z.umn.edu/etog-etrg
@@ I Introduction
Recent robot grasping studies using Large Language Models (LLMs) or Multimodal Large Language Models (MLLMs) have demonstrated strong object grounding and grasp detection performance [16, 17, 18, 19, 20]. However, full-model training or fine-tuning is computationally expensive and may suffer from catastrophic forgetting [21], limiting the feasibility of local deployment. In this paper, we propose a parameter-efficient tuning (PET) approach [22, 23] for object grounding and robot grasping, enabling competitive performance by incorporate a lightweight module into a frozen pre-trained model.
@@ IV Experiments
TABLE I: Performance comparison of different SOTA CLIP-based models on RefCOCO related datasets using mIoU metric. Backbone: visual encoder architecture. U: the UMD test split for RefCOCO(g). Params: total number of parameters to be tuned.
@@
ETOG (Ours) ICRA25 CLIP-R50 72.31 75.49 66.62 63.26 69.46 53.27 60.80 61.98 65.40 32.93M
@@
ETOG (Ours) ICRA25 CLIP-R101 73.37 76.16 68.54 64.70 70.86 54.94 63.06 63.66 66.91 30.57M
@@
ETOG (Ours) ICRA25 CLIP-ViT-B 73.37 76.90 69.34 65.95 71.47 56.94 63.75 64.63 67.79 29.89M
@@
ETOG (Ours) ICRA25 ResNet50 70.11 73.97 63.76 58.63 65.66 47.86 57.35 59.87 62.15
ETOG (Ours) ICRA25 ResNet101 71.11 74.30 65.61 60.00 67.27 49.73 60.09 61.23 63.67
ETOG (Ours) ICRA25 ViT-B 71.35 76.14 66.73 62.33 68.52 51.90 61.12 62.85 65.12
@@ IV-A Referring Expression Segmentation for Object Grounding
ETOG Comparison with RES Baselines: In Table II, we compare our ETOG with the state-of-the-art CLIP-based RES methods on RefCOCO datasets using mIoU metric. ETOG outperforms all baseline methods across different backbone architectures. CLIPORT-semantic, a variant of CLIPORT [31], excludes the depth processing branch and uses simple element-wise multiplication to fuse vision and language features. The comparison between CLIPORT-semantic and our method highlights the effectiveness of our VL-adapter. In particular, with the CLIP [1] ResNet-101 [38] backbone, our method–using only $1.9$M tunable parameters–achieves a $2.66\%$ and $1.95\%$ performance gain across 8 evaluation tasks compared to CRIS [26] and EAVL [27] that are fully CLIP fine-tuned methods. Additionally, ETOG with $1.21$M tunable ViT-B-16 backbone parameters shows a $4.35\%$ and $0.73\%$ performance improvement compared to the latest parameter-efficient tuning approaches, ETRIS [22] and BARLERIA [23], respectively. Table II further compares our model with full fine-tuned RES models, showing that ETOG achieves competitive object grounding performance even against non-CLIP-based approaches. Fig. 3 provides qualitative results, including model predictions and feature map visualizations, with the fourth column showing the attention regions of the well-fused multimodal feature $f_{c}$.
@@ IV-B Referring Grasp Synthesis with Depth Fusion
Ablation Study on Depth Fusion Branch: One primary SOTA baseline we evaluate against is CROG [11], a CLIP-based RGS method with full model fine-tuned with OCID-VLG dataset. As shown in Table IV, ETRG-A without depth model (ResNet-50) achieves a slightly better ($+3.96\%$ $J@1$) performance and competitive object grounding (RES) performance compared CROG. After integrating depth inputs into VL-adapter (details in Fig. 2) fusing the object geometric information into visual-language branches, our ETRG-A with depth fusion full model attains higher grasping rectangle prediction ($+11.86\%$ $J@1$) accuracy with ResNet-50 backbone and ($+12.18\%$ $J@1$) accuracy with ResNet-101 backbone compared with CROG. CROG is marginally better than our models in the RES task, but we attribute this to the simplicity of the OCID-VLG dataset, which primarily features tabletop objects and lacks the scene diversity found in RefCOCO-related datasets. This environment favors full model fine-tuning methods like CROG, potentially leading to over-fitting. Although CROG has 147M tunable parameters which is 3 times larger than our proposed depth-fusion approach, our ETRA-A model still achieves competitive performance on this RES task.
@@ IV-C Referring Grasp Affordance for Spatial Reasoning
ETRG-B (Ours) 90.75 88.38 86.56 88.56
Simulation Results: We evaluate our methods on affordance prediction with spatial reasoning. The experiments include two primary scene types: Absolute and Relative. In each test scene, there are three identical object instances, and spatial language expression are used to specify the target. To increase the difficulty of spatial reasoning, we create test cases with four objects and seven objects. As illustrated in Table V, both the proposed methods, with and without depth fusion, demonstrate strong performance in handling complex spatial reasoning tasks. Our model with depth fusion achieves slightly higher performance compared to the RGB-only method.
@@
Real Robot Results: We conducted experiments using a Franka Emika Panda robot equipped with a FESTO DHAS soft gripper. We collected over 10 household objects and created scenes with 6 randomly selected objects. Each scene included up to 3 duplicated instances of one category to test spatial reasoning. For each baseline, we report object grounding accuracy, which assesses whether the grasp is on the correct target, and the overall grasp success rate. A detailed visualization of correct mustard bottle identification with feasible grasp poses is shown in Fig. 4. Across 40 grasp attempts with different object arrangements (Table VI), both methods achieve comparable grounding performance, with our model incorporating depth information showing slightly better grasping accuracy.
TABLE VI: ETRG-B Real robot results on RGA spatial reasoning.
@@ V Conclusion
In this work, we presented a novel CLIP-based parameter-efficient tuning framework that can be adapted to a range of tasks, including RES, RGS, and RGA tasks. Our method incorporated a vision-language fusion adapter and a depth-fusion branch, enabling effective open-vocabulary object grounding and robot grasping. We further explored the spatial reasoning capabilities in robot grasping, which the original CLIP [1] model does not address. The proposed models demonstrated competitive performance across these challenging vision language and robotics tasks. The current method may under-perform in highly cluttered scene.